OpenStudy (waheguru):

I need to convert this to standard form y=-1/3x+5/6

Parth (parthkohli):

Just take the \(x\) along with its coefficient to the Left-Hand-Side.

Parth (parthkohli):

Since the standard form is \(Ax + By = C\), x and y stay together.

OpenStudy (waheguru):

Is it 2x-5y-30

Parth (parthkohli):

Nope.

Parth (parthkohli):

The only work to be done here is adding 1x/3 to both sides.

OpenStudy (waheguru):

y=-1/3x+5/6 now I times 6 6y=-6/3+30/6 right

Parth (parthkohli):

What?

Parth (parthkohli):

No need to complicate it!

Parth (parthkohli):

Just add -1x/3 to both sides, and you get the standard form.

OpenStudy (waheguru):

what?

OpenStudy (waheguru):

Woops, I comfused the answer with the qrong question the answer was 2x+6y-5=0

Parth (parthkohli):

\[y + {1x \over 3} =\cancel{ {-1 \over 3x} + {1x \over 3} }+ {5 \over 6}\]\[ \color{Black}{\Rightarrow y + {x \over 3} = {5 \over 6}}\]

Parth (parthkohli):

Huh? What question's answer was that?

OpenStudy (waheguru):

For y=2/5x-3 is 2x-5y-15

OpenStudy (precal):

looks like the multiplied it by 6 to get rid of all of the fractions

OpenStudy (precal):